OpenStudy (anonymous):

3. Write an explicit formula for the sequence1/2,3/7,1/3,5/19,3/14 , ... Then find a(14). (1 point)

OpenStudy (anonymous):

15/99 i belive is a(14) but i need help writing the formula

OpenStudy (anonymous):

15/199*

OpenStudy (amistre64):

1/2, 3/7, 1/3, 5/19, 3/14 hmm, is the average of 1/2 and 1/3, 3/7 by any chance?

OpenStudy (amistre64):

well, its not arithmatic then .... 1/2 * (2*3/7) = 3/7 is 3/7 * (2*3/7) = 1/3 ? nope,, so it aint geometirc

OpenStudy (amistre64):

1/2, 3/7, 1/3, 5/19, 3/14 (1, 6/7, 2/3, 10/19, 3/7) /2 (7, 6, 14/3, 70/19, 3) /14 (21, 18, 14, 210/19, 9) /42 (399, 342, 266, 210, 171) /798 that at least gets us all common denominator if ituseful

OpenStudy (amistre64):

why do you think 15/99 is it?